P can do a piece of work in 20 days and Q in 24 days. find how much time will they take to complete the work if they work alternately starting with P ??
Answer by richwmiller(17219)

1/a+1/b=1/x
1/20+1/24=1/x
24/480+20/480=1/x
44/480=1/x
x=480/44
x=10.9090909 days working together
Double the time if they don't work at the same time